Barry is 8 years older than his sister. In 3 years, he will be twice as old as she will be then. How old is each now?
Answer by Maths68(1474)

Let
Present age of Barry = b
Present age of his sister = s


Barry is 8 years older than his sister
b=s+8..........(1)


In 3 years
Barry will be = b+3
His sister will be = s+3
Then
He will be twice as old as his sister will be
b+3=2(s+3)
b+3=2s+6
b=2s+6-3
b=2s+3.................(2)
Substitute the value of b from (1) to (2)
s+8=2s+3
s-2s=3-8
-s=-5
cancel out (-)
s=5

Substitute the value of s in (1)
b=s+8
b=5+8
b=13

Present age of Barry = b = 13 years old
Present age of his sister = s = 5 years old